Solve, rounding to the nearest tenth:
On a sunny day, a tree and its shadow form the sides of a right triangle. If the hypotenuse is 52 m long and the tree is 48 m tall, how long is the shadow?

Answer by checkley77(12844)

48^2+X^2=52^2
2304+X^2=2704
X^2=2704-2304
X^2=400
X=SQRT400
X=20 FEET LONG FOR THE SHADOW.